sum of a number and its double is 9.Find the number
Found 2 solutions by Cromlix, lenny460:
Answer by Cromlix(4381)

x + 2x = 9
3x = 9
x = 3
Your answer is 3

Answer by lenny460(1073)

Sum of a number and its double is 9. Find the number ?

3 + 6 = 9
The number is 3.
It's double is 3 * 2 = 6
So the sum of a number and its double is 9.


The number is: 3

Answer:
3





Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com